# Beneath the Surface: How Large Language Models Reflect Hidden Bias

### v1

@@ 6 Conclusion
In this work, we propose the Hidden Bias Benchmark (HBB), a novel dataset designed to systematically assess hidden bias in LLMs. Unlike previous benchmarks that focus on overt bias through direct demographic term associations, HBB evaluates how biases persist in real-world narratives where stereotypes are contextually hidden rather than explicitly stated. We detail HBB’s construction, demonstrating how bias concepts and demographic descriptors are subtly hidden into realistic scenarios. To rigorously evaluate hidden bias, we measure response variations across parallel test instances. And we conduct an extensive analysis to examine how biases manifest across different models, demographic categories, identities, and descriptors. Our findings reveal that while LLMs exhibit reduced bias in response to overt bias, they continue to reinforce bias in subtle, hidden contexts. This highlights HBB’s value as a complementary tool for bias measurement, addressing limitations of previous benchmarks.
@@
Comparability between HBB and HBB-SG
Our HBB-SG (semi-generation) analysis cannot be directly compared to HBB (MCQ-based evaluation) due to fundamental differences in evaluation metrics. MCQ settings constrain models to predefined answer options, whereas semi-generation measures models’ generated responses based on perplexity and converts them into probability scores later, making biases harder to quantify in a directly comparable manner. Future work should refine methodologies for aligning results across these evaluation paradigms. Intuitively, generation-based models may exhibit greater bias in free-form text compared to multiple-choice settings. In real-world applications, LLMs do not operate under rigid MCQ structures but instead generate open-ended responses, where biases may be more pronounced. Future studies should further investigate how bias manifests in long-form generation to better reflect real-world usage.
@@
Currently, HBB evaluates bias across five social categories (Age, Race Ethnicity, Gender, Socioeconomic Class, and Religions). However, many other demographic categories, such as disability status or physical appearance, remain unexplored. Expanding the dataset to incorporate a broader range of identities would enable a more comprehensive fairness assessment.
@@
HBB currently derives its bias concepts from well-known bias benchmarks such as BBQ, SOFA, CrowS-Pairs, and StereoSet. While these datasets provide a strong foundation, they may not fully capture all real-world biases. Future iterations of HBB should incorporate more diverse, dynamically generated biases, leveraging data-driven stereotype discovery methods to enrich the dataset with emerging and underrepresented biases.
@@
Our dataset is adaptable to any language, our experiments focus on English due to the scarcity of annotated stereotype datasets in other languages. We strongly advocate for the creation of multilingual datasets to facilitate bias assessment in LLMs, as demonstrated in Martinková et al. (2023); Zhao et al. (2024); Fleisig et al. (2024).
@@
Our bias analysis requires evaluating each question ten times to estimate answer probabilities, making it both computationally expensive given current OpenAI API pricing and inefficient. Moreover, analyzing all test instances further reduces efficiency. Future research could optimize this process by leveraging output token probabilities to approximate answer selections and concentrating on test instances ($\geq 20$ bias score) identified in HBB for bias analysis.
@@
HBB is designed to assess hidden biases in LLMs by systematically hidden bias-related concepts within subtly framed contexts. HBB extracts bias concepts exclusively from well-established bias evaluation datasets, including CS, SS, BBQ, and SOFA, ensuring that all stereotypes and demographic categories originate from prior research. Our benchmark focuses on five demographic categories – Age, Gender, Race Ethnicity, Socioeconomic Class, and Religions – providing a structured but non-exhaustive examination of social biases. While these categories cover a range of biases, they do not comprehensively capture the full complexity of demographic identities.
HBB does not introduce new bias concepts; rather, it relies on existing datasets that may already contain biases inherent in their original sources, such as Western societal norms. As bias perception is highly context-dependent, our benchmark may not fully account for intersectional biases or regional and cultural variations in stereotype formation. Additionally, while HBB evaluates biases by comparing responses across demographic descriptors, reducing bias assessment to a single metric has inherent limitations. Bias manifests in complex ways that cannot always be fully captured through automated benchmarks alone.
Thus, we advocate for the responsible use of our HBB, emphasizing that it should serve as a complementary tool rather than a definitive measure of bias. Researchers and practitioners are encouraged to use HBB alongside qualitative human analysis, and to refine and expand the dataset to enhance its inclusivity and applicability across broader social contexts.
@@ Appendix A Model Size and Computational Budget
We utilize six recent LLMs: GPT-4o (gpt-4o-20240513) Hurst et al. (2024), Llama-3.2-11B-Vision-Instruct, Llama-3.2-3B-Instruct, and Llama-3.1-8B-Instruct Dubey et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), and Qwen2.5-7B-Instruct Team (2024). For our experiments, we set temperature = 0.8, top_p = 1, frequency_penalty = 0.6, no presence penalty, no stopping condition other than the maximum number of tokens to generate, max_tokens = 2048. All experiments are conducted on AMD - 1984 cores CPUs and an Nvidia A100 - 80GB GPUs. For our HBB, It takes less than 30 minutes for GPT-4o Batch API to evaluate all questions. Llama-3.2-11B-Vision-Instruct needs around 21 hours to run all questions in our HBB. Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct take approximately 18 hours to run all questions in HBB. And Llama-3.2-3B-Instruct finishes all questions in HBB less than 10 hours.
@@ Appendix B Related Work
Overt Bias Benchmarks.
@@
Overt bias in LLMs has been widely examined using benchmarks that evaluate whether LLMs systematically favor stereotypical responses over anti-stereotypical ones when provided with explicit demographic identities. And multiple benchmarks have been designed to quantify overt bias from diverse perspectives, facilitating structured evaluations of LLM bias Parrish et al. (2022); Nangia et al. (2020); Nadeem et al. (2021); Marchiori Manerba et al. (2024); Bi et al. (2023); del Arco et al. (2024); Kotek et al. (2023).
@@
CrowS-Pairs (CS) Nangia et al. (2020) and StereoSet (SS) Nadeem et al. (2021) are among the first benchmarks designed to systematically evaluate social biases in LLMs. CS features sentence pairs, one containing a stereotypical statement and the other presenting an anti-stereotypical alternative. Log-likelihood comparisons reveal whether models systematically favor stereotypical associations. SS extends this approach to both masked and autoregressive LMs, computing a stereotype score that quantifies model preference for stereotypical completions over neutral alternatives. BBQ Parrish et al. (2022) enhances explicit bias evaluation by incorporating ambiguous and disambiguated question formats to analyze bias in structured reasoning tasks to assess whether models rely on stereotypes in QA tasks, distinguishing responses with and without informative context to reveal how bias affects decision-making. And SOFA Marchiori Manerba et al. (2024) extends bias evaluation by incorporating a broader range of stereotypes and demographic identities, moving beyond binary group comparisons. Together, these benchmarks establish the foundation for overt bias evaluation, assessing how LLMs respond to overtly biased statements.
Hidden Bias and the Evolution of Model Behavior.
@@
As LLMs advance, their responses to overt bias evaluations have become more neutral and self-regulated, often producing answers that align with socially desirable norms. This shift is largely due to improvements in model training, particularly through methods such as instruction tuning and alignment techniques that reinforce neutrality in responses to explicitly biased contexts Ouyang et al. (2022); Zhang et al. (2023); Peng et al. (2023); Ji et al. (2024). Consequently, traditional overt bias benchmarks mentioned previously, often show reduced bias scores for LLMs. However, the absence of overt bias in model responses does not necessarily indicate genuine bias mitigation; rather, biases may persist in subtler, more hidden ways that traditional evaluation methods fail to capture.
@@
Recent studies Bai et al. (2024); Smith et al. (2022) have sought to evaluate implicit bias in LLMs by expanding beyond direct stereotype statements. Bai et al. (2024) measure bias by prompting LLMs to associate specific words with demographic identities and subsequently using these associations to generate narratives. This approach seeks to identify decision-making biases by examining whether LLMs systematically generate narratives that reinforce demographic stereotypes. Nonetheless, the method relies on manually selecting words associated with demographics, which does not necessarily reflect biased model behavior but rather linguistic appropriateness in describing certain groups. LLMs may generate contextually appropriate descriptions rather than perpetuating harmful stereotypes. Our proposed Hidden Bias Benchmark (HBB) diverges from this approach by hiding biases within naturally occurring contextual narratives instead of relying on predefined word associations. Unlike Bai et al. (2024), which explicitly links demographic identities to predefined words, HBB examines biases when bias concepts are subtly hidden in realistic scenarios. Rather than prompting models to associate words with demographic groups, HBB evaluates bias by analyzing response variations across parallel test instances with the same demographic descriptor type, where biases are subtly hidden in naturalistic language.
@@ Appendix C Dataset Generation
C.1 Pairs of Opposite Bias Related Concepts
@@ C.1.1 Bias Summary
Table 12 shows all prompts for each dataset to generate bias summary.
@@ C.1.2 Raw Concept Pairs
We also provide several examples after the final answers format in the prompt to conduct in-context learning Brown et al. (2020) in order to retrieve high quality concept pairs. The basic prompt structures are in Table 13.
@@ C.3 Data Construction
Table 10 summarizes all subtle replacements for various identities, while Table 4 lists all names used to replace [[X]]. And Table 5 shows statistics of each category in HBB.
@@ D.1 Metrics for Baseline Datasets
Furthermore, regarding Section 4.2.2, we utilize bias measurements from each dataset baseline to compare the severity of bias across different baseline models. Specifically, we conduct MCQ bias evaluation for our dataset. For BBQ-ambig, we use the ambiguous bias score Parrish et al. (2022) with range of (-1, 1) and 0 indicates no bias. For BBQ-disambig, we directly compute the accuracy of correct answers, as it serves as the most reliable indicator for disambiguated text, which ranges from 0 to 100, where 0 demonstrates highest bias and 100 shows no bias. We apply the probability bias score from Nangia et al. (2020) for the CS dataset, where a score of 50 indicates neutrality with no bias within the range of (0, 100). Moreover, we utilize the ICAT score Nadeem et al. (2021) to measure bias levels in SS datasets. In this scoring system, which ranges from 0 to 100, a score of 0 represents the most severe bias, while 100 indicates no bias. We use prompt in Table 11 for LLMs to evaluate bias.
@@
Figure 4: Hidden bias score distributions for HBB.
D.2 Bias Analysis in HBB
HBB reveals biases across different models, with GPT-4o exhibiting the highest bias score. The first two columns in Table 1 present the average bias score and total count of all test instances ($\geq 20$ bias score), indicating that every model exhibits some degree of social bias. And Figure 4 shows bias score distributions across models.
Model HBB BBQ-ambig BBQ-disambig CS SC-intra SC-inter
@@ D.2 Bias Analysis in HBB
Figure 5: Bias score distribution of sample from HBB for GPT-4o (corresponding test instances from BBQ-ambig).
@@
Table 7: Top bias associations for each category in HBB (GPT-4o).
Figure 6: HBB reflects bias (correspond to same bias concept from BBQ) 1.
Figure 7: HBB reflects bias (correspond to same bias concept from BBQ) 2.
@@ D.3 Bias Analysis cross datasets
More advanced models show higher hidden bias but lower overt bias, whereas less advanced models display the opposite trend. In addition to bias scores for measuring bias, we assess the refuse rate as an indicator of both model comprehension and dataset quality, as shown in Table 6, to provide further insight into bias scores. The refuse rate refers to the proportion of questions where the model either fails to follow the instructions in the prompt (Table 11 in Appendix) or declines to answer. GPT-4o demonstrates superior comprehension and response effectiveness compared to other models, and HBB maintains high quality for questions, as evidenced by models’ willingness to generate responses.
For the same bias concepts, LLMs exhibit bias in HBB, but show no bias in previous datasets. The distribution of test instances is shown in Figure 5. Refers to Figure 6 and Figure 7 as additional examples for the corresponding BBQ bias concept and our HBB test instance. These findings suggest that HBB detects substantially higher bias for the same concepts, demonstrating that LLMs still exhibit nuanced biases closely mirroring real-world scenarios.
HBB can be used to discover bias. Table 7 presents top test instances with 100 bias score, and show bias related concept pairs associated with specific demographic identities for each category.
D.3.1 Discussion
It is important to note that although the CrowS-Pairs (CS) dataset exhibits relatively higher bias scores, the dataset contains numerous questions of poor quality. Blodgett et al. (2021) highlights that many examples in the CS dataset do not effectively study biases, and the design of numerous biased answer options is often confusing. Specifically, the study found that many benchmark datasets used for assessing bias in language models suffer from validity issues. In particular, the contrastive sentence pairs in CS often lack clear conceptualization and operationalization of stereotypes, which undermines the reliability of bias evaluations. As a result, the high bias scores observed in these previous s should be interpreted with caution, as they may be influenced by the dataset’s inherent design flaws rather than genuine model biases. Our proposed HBB, which features well-defined answer options and more realistic scenario descriptions for each question, provides a more effective design for identifying bias.
Appendix E Semi-Generation Based HBB (HBB-SG)
E.1 HBB-SG Bias Measures
Based on the same bias measurement mechanism in Section 3.3, the probability of selecting an answer option for Question 1 option A, for example,$P_{1}(A)$, is computed as the average reciprocal of perplexity (PPL) Jelinek et al. (1977) across all generated variations:

### v2

@@ 6 Conclusion
In this work, we introduce the Description-based Bias Benchmark (DBB), a novel dataset for systematically evaluating bias by the description-based method in LLMs. Unlike prior benchmarks that assess bias via explicit demographic-term associations to form term-based evaluation, DBB captures how biases persist in realistic depictions where stereotypes are subtly hidden. We detail DBB’s construction, where demographic descriptors and bias concepts are hidden within naturalistic contexts, and evaluate model responses across parallel test instances. Our analysis reveals that while LLMs show reduced bias in term-based evaluation, they continue to reinforce bias in subtle, descriptive settings. This highlights DBB’s value as a complementary tool for bias measurement, addressing the limitations of previous benchmarks.
Acknowledgments
This work is in part supported by NSF grant IIS-2452129 and the Commonwealth Cyber Initiative (CCI) grant (HN-4Q24-055). Some LLM resources were accessed from the Accelerating Foundation Models Research program at Microsoft Research awarded to ZY.
@@ Acknowledgments
Comparability between DBB and DBB-SG
Our DBB-SG (semi-generation) analysis cannot be directly compared to DBB (MCQ-based evaluation) due to fundamental differences in evaluation metrics. MCQ settings constrain models to predefined answer options, whereas semi-generation measures models’ generated responses based on perplexity and converts them into probability scores later, making biases harder to quantify in a directly comparable manner. Future work should refine methodologies for aligning results across these evaluation paradigms. Intuitively, generation-based models may exhibit greater bias in free-form text compared to multiple-choice settings. In real-world applications, LLMs do not operate under rigid MCQ structures but instead generate open-ended responses, where biases may be more pronounced. Future studies should further investigate how bias manifests in long-form generation to better reflect real-world usage.
@@
Currently, DBB evaluates bias across five social categories (Age, Race Ethnicity, Gender, Socioeconomic Class, and Religions), using descriptors adapted from and inspired by prior studies such as BBQ, SOFA, CrowS-Pairs, and StereoSet. However, many other demographic categories, such as disability status or physical appearance, remain unexplored. In addition, the current set of descriptors may not fully capture the diversity within each category. Expanding the dataset to incorporate a broader range of identities and richer descriptors would enable a more comprehensive fairness assessment.
@@
DBB currently derives its bias concepts from well-known bias benchmarks such as BBQ, SOFA, CrowS-Pairs, and StereoSet. While these datasets provide a strong foundation, they may not fully capture all real-world biases. Future iterations of DBB should incorporate more diverse, dynamically generated biases, leveraging data-driven stereotype discovery methods to enrich the dataset with emerging and underrepresented biases.
@@
Our dataset is adaptable to any language; our experiments focus on English due to the scarcity of annotated stereotype datasets in other languages. We strongly advocate for the creation of multilingual datasets to facilitate bias assessment in LLMs, as demonstrated in Martinková et al. (2023); Zhao et al. (2024a); Fleisig et al. (2024).
@@
Our bias analysis requires evaluating each question ten times to estimate answer probabilities, making it both computationally expensive given current OpenAI API pricing and inefficient. Moreover, analyzing all test instances further reduces efficiency. Future research could optimize this process by leveraging output token probabilities to approximate answer selections and concentrating on test instances ($\geq 20$ bias score) identified in DBB for bias analysis.
Automatic Qualitative Evaluation
Our DBB lacks an automatic qualitative evaluation to systematically verify whether all the contexts and options naturally reflect the intended bias concepts. While we manually ensure coherence and semantic alignment during data construction, future work could explore automated methods to assess contextual relevance and concept clarity at scale.
@@
DBB is designed to assess biases in LLMs by a systematic description-based approach. DBB extracts bias concepts exclusively from well-established bias evaluation datasets, including CS, SS, BBQ, and SOFA, ensuring that all stereotypes and demographic categories originate from prior research. Our benchmark focuses on five demographic categories – Age, Gender, Race Ethnicity, Socioeconomic Class, and Religions – providing a structured but non-exhaustive examination of social biases. While these categories cover a range of biases, they do not comprehensively capture the full complexity of demographic identities.
DBB does not introduce new bias concepts; rather, it relies on existing datasets that may already contain biases inherent in their original sources, such as Western societal norms. As bias perception is highly context-dependent, our benchmark may not fully account for intersectional biases or regional and cultural variations in stereotype formation. Additionally, while DBB evaluates biases by comparing responses across demographic descriptors, reducing bias assessment to a single metric has inherent limitations. Bias manifests in complex ways that cannot always be fully captured through automated benchmarks alone.
Thus, we advocate for the responsible use of our DBB, emphasizing that it should serve as a complementary tool rather than a definitive measure of bias. Researchers are encouraged to use DBB alongside qualitative human analysis, and to refine and expand the dataset to enhance its inclusivity and applicability across broader social contexts.
@@ Appendix A Model Size and Computational Budget
We utilize six recent LLMs: GPT-4o (gpt-4o-20240513) Hurst et al. (2024), Llama-3.2-11B-Vision-Instruct, Llama-3.2-3B-Instruct, and Llama-3.1-8B-Instruct Dubey et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), and Qwen2.5-7B-Instruct Team (2024). For our experiments, we set temperature = 0.8, top_p = 1, frequency_penalty = 0.6, no presence penalty, no stopping condition other than the maximum number of tokens to generate, max_tokens = 2048. All experiments are conducted on AMD - 1984 cores CPUs and Nvidia A100 - 80GB GPUs. For our DBB, it takes less than 30 minutes for the GPT-4o Batch API to evaluate all questions. Llama-3.2-11B-Vision-Instruct needs around 21 hours to run all questions in our DBB. Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct take approximately 18 hours to run all questions in DBB. And Llama-3.2-3B-Instruct finishes all questions from DBB in less than 10 hours.
@@ Appendix B Related Work
Term-Based Evaluation
@@
Social bias Zhao et al. (2024b); Venkit et al. (2022); Field and Tsvetkov (2020); Lin and Li (2025); Tan and Lee (2025) in LLMs has been widely examined using benchmarks that evaluate whether LLMs systematically favor stereotypical terms over anti-stereotypical ones when provided with explicit demographic identities. And multiple benchmarks have been designed to quantify bias at the term level from diverse perspectives, facilitating structured evaluations of LLM bias Parrish et al. (2022); Nangia et al. (2020); Nadeem et al. (2021); Marchiori Manerba et al. (2024); Bi et al. (2023); del Arco et al. (2024); Kotek et al. (2023).
@@
CrowS-Pairs (CS) Nangia et al. (2020) and StereoSet (SS) Nadeem et al. (2021) are among the first benchmarks designed to systematically evaluate social biases in LLMs. CS features sentence pairs, one containing a stereotypical statement and the other presenting an anti-stereotypical alternative. Log-likelihood comparisons reveal whether models systematically favor stereotypical associations. SS extends this approach to both masked and autoregressive LMs, computing a stereotype score that quantifies model preference for stereotypical completions over neutral alternatives. BBQ Parrish et al. (2022) enhances explicit bias evaluation by incorporating ambiguous and disambiguated question formats to analyze bias in structured reasoning tasks to assess whether models rely on stereotypes in QA tasks, distinguishing responses with and without informative context to reveal how bias affects decision-making. SOFA Marchiori Manerba et al. (2024) extends bias evaluation by incorporating a broader range of stereotypes and demographic identities, moving beyond binary group comparisons. SEAT May et al. (2019) and WEAT Caliskan et al. (2017) evaluate social bias by measuring associations in the embedding space using sentence or word encoders. However, these methods still rely on explicit demographic and attribute terms, making them a form of term-based evaluation. And BOLD dataset Dhamala et al. (2021) prompts models with identity-specific sentences and analyzes generated continuations for polarity and toxicity across domains such as gender, religion, and profession. Despite using generative outputs, BOLD, and similar methods still operate under the term-based paradigm, as they rely on superficial demographic identifiers in prompts to elicit biased behavior. Together, these benchmarks establish the foundation for term-based bias evaluation, assessing how LLMs respond to superficially biased statements.
Description-Based Evaluation
@@
As LLMs advance, their responses to term-based bias evaluations have become more neutral and self-regulated, often producing answers that align with socially desirable norms. This shift is largely due to improvements in model training, particularly through methods such as instruction tuning and alignment techniques that reinforce neutrality in responses to explicitly biased contexts Ouyang et al. (2022); Zhang et al. (2023); Peng et al. (2023); Ji et al. (2024). Consequently, the traditional term-based bias benchmarks mentioned previously, often show reduced bias scores for LLMs. However, the absence of bias under term-based evaluation in model responses does not necessarily indicate genuine bias mitigation; rather, biases may persist in subtler, more hidden ways that traditional term-based bias evaluation methods fail to capture.
Therefore, social bias measurement can be divided into two modes: term-based evaluation and description-based evaluation. Term-based bias evaluation measures associations between demographic identities and explicit stereotype terms – typically at the lexical level. In contrast, description-based bias evaluation evaluates associations between demographic identities and bias-related concepts hidden in naturalistic descriptions, often conveyed through behaviors, traits, and more – capturing bias at the semantic level. Both manifest the same underlying bias but differ in the form: one through direct demographic-term associations, the other through contextualized demographic-concept associations. Existing benchmarks are all term-based evaluations. In this work, our bias definition aligns with representational harms by Blodgett et al.; Gonçalves and Strubell; Crawford – Differences in system performance for different social groups, language that misrepresents the distribution of different social groups in the population, or language that is denigrating to particular social groups.
@@
Though recent studies Bai et al. (2024); Smith et al. (2022) have sought to evaluate bias in LLMs by expanding beyond direct stereotype statements. Bai et al. (2024) measures bias by prompting LLMs to associate specific words with demographic identities and subsequently using these associations to generate narratives. This approach seeks to identify decision-making biases by examining whether LLMs systematically generate narratives that reinforce demographic stereotypes. Nonetheless, the method relies on manually selecting words associated with demographics, which does not necessarily reflect biased model behavior but rather linguistic appropriateness in describing certain groups. LLMs may generate contextually appropriate descriptions rather than perpetuating harmful stereotypes. Although this approach represents a more nuanced application of term-based bias evaluation, the method still relies on explicitly linking demographics to predefined word associations (still under term-based mode). Our proposed Description-based Bias Benchmark (DBB) diverges from this approach by hiding biases within naturally occurring contextual narratives instead of relying on predefined word associations. Unlike Bai et al. (2024), which explicitly links demographic identities to predefined words, DBB examines biases when bias concepts are subtly hidden in realistic scenarios. Rather than prompting models to associate words with demographic groups, DBB evaluates bias by analyzing response variations across parallel test instances with the same demographic descriptor type, where biases are subtly hidden in naturalistic language.
@@ Appendix C Dataset Generation
C.1 Pairs of Opposite Bias-Related Concepts
@@ C.1.1 Bias Summary
Table 12 shows all prompts for each dataset to generate the bias summary.
@@ C.1.2 Raw Concept Pairs
We also provide several examples after the final answers format in the prompt to conduct in-context learning Brown et al. (2020) in order to retrieve high-quality concept pairs. The basic prompt structures are in Table 13.
@@ C.3 Data Construction
Table 10 summarizes all subtle replacements for various identities, while Table 4 lists all names used to replace [[X]]. And Table 5 shows statistics of each category in DBB.
@@ D.1 Metrics for Baseline Datasets
Furthermore, regarding Section 4.2.2, we utilize bias measurements from each dataset baseline to compare the severity of bias across different baseline models. Specifically, we conduct MCQ bias evaluation for our dataset. For BBQ-ambig, we use the ambiguous bias score Parrish et al. (2022) with a range of (-1, 1), and 0 indicates no bias. For BBQ-disambig, we directly compute the accuracy of correct answers, as it serves as the most reliable indicator for disambiguated text, which ranges from 0 to 100, where 0 demonstrates the highest bias and 100 shows no bias. We apply the probability bias score from Nangia et al. (2020) for the CS dataset, where a score of 50 indicates neutrality with no bias within the range of (0, 100). Moreover, we utilize the ICAT score Nadeem et al. (2021) to measure bias levels in SS datasets. In this scoring system, which ranges from 0 to 100, a score of 0 represents the most severe bias, while 100 indicates no bias. We use the prompt in Table 11 for LLMs to evaluate bias.
@@
Figure 4: Bias score distributions for DBB.
D.2 Bias Analysis in DBB
DBB reveals biases across different models, with GPT-4o exhibiting the highest bias score. The first two columns in Table 1 present the average bias score and total count of all test instances ($\geq 20$ bias score), indicating that every model exhibits some degree of social bias. And Figure 4 shows bias score distributions across models.
Model DBB BBQ-ambig BBQ-disambig CS SC-intra SC-inter
@@ D.2 Bias Analysis in DBB
Figure 5: Bias score distribution of sample from DBB for GPT-4o (corresponding test instances from BBQ-ambig).
@@
Table 7: Top bias associations for each category in DBB (GPT-4o).
Figure 6: DBB reflects bias (corresponds to the same bias concept from BBQ) 1.
Figure 7: DBB reflects bias (corresponds to the same bias concept from BBQ) 2.
@@ D.3 Bias Analysis cross datasets
More advanced models show a higher level of bias in description-based evaluation but a lower level of bias via the term-based method, whereas less advanced models display the opposite trend. In addition to bias scores for measuring bias, we assess the refuse rate as an indicator of both model comprehension and dataset quality, as shown in Table 6, to provide further insight into bias scores. The refuse rate refers to the proportion of questions where the model either fails to follow the instructions in the prompt (Table 11 in Appendix) or declines to answer. GPT-4o demonstrates superior comprehension and response effectiveness compared to other models, and DBB maintains high quality for questions, as evidenced by the models’ willingness to generate responses.
DBB can be used to discover bias. Table 7 presents top test instances with a bias score of 100, and shows bias-related concept pairs associated with specific demographic identities for each category.
D.4 Instance Match: DBB vs. BBQ
For the same bias concepts, LLMs exhibit bias in DBB, but show no bias in previous datasets. The distribution of test instances is shown in Figure 5. Refers to Figure 6 and Figure 7 as additional examples for the corresponding BBQ bias concept and our DBB test instance. These findings suggest that DBB detects substantially higher bias for the same concepts, demonstrating that LLMs still exhibit nuanced biases closely mirroring real-world scenarios.
D.4.1 Discussion
It is important to note that although the CrowS-Pairs (CS) dataset exhibits relatively higher bias scores, the dataset contains numerous questions of poor quality. Blodgett et al. (2021) highlights that many examples in the CS dataset do not effectively study biases, and the design of numerous biased answer options is often confusing. Specifically, the study found that many benchmark datasets used for assessing bias in language models suffer from validity issues. In particular, the contrastive sentence pairs in CS often lack clear conceptualization and operationalization of stereotypes, which undermines the reliability of bias evaluations. As a result, the high bias scores observed in these previous studies should be interpreted with caution, as they may be influenced by the dataset’s inherent design flaws rather than genuine model biases. Our proposed DBB, which features well-defined answer options and more realistic scenario descriptions for each question, provides a more effective design for identifying bias.
Appendix E Semi-Generation Based DBB (DBB-SG)
E.1 DBB-SG Bias Measures
[…]